1.    Assuming that the samples come from normal distributions, find the margin of error  given the following:

a. n = 10 and X = 28 with s = 4.0, 90% confidence

b. n = 16 and X = 50 with s = 4.2, 95% confidence

c. n = 20 and X = 68.2 with s = 2.5, 90% confidence

d. n = 23 and X = 80.6 with s = 3.2, 95% confidence

e. n = 25 and X = 92.8 with s = 2.6, 99% confidence

 

2.    Using the information in number 2, find the interval estimates of the population mean.


1
Expert's answer
2022-06-07T05:50:52-0400

1.

a.

The critical value for "\alpha = 0.10, df=n-1=9" degrees of freedom is "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}=1.833113."

"E=t_c\times\dfrac{s}{\sqrt{n}}=1.833113\times\dfrac{4}{\sqrt{10}}=2.3187"

b.

The critical value for "\alpha = 0.05, df=n-1=15" degrees of freedom is "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}=2.131449."

"E=t_c\times\dfrac{s}{\sqrt{n}}=2.131449\times\dfrac{4.2}{\sqrt{16}}=2.2380"



c.

The critical value for "\alpha = 0.10, df=n-1=19" degrees of freedom is "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}=1.729133."

"E=t_c\times\dfrac{s}{\sqrt{n}}=1.729133\times\dfrac{2.5}{\sqrt{20}}=0.9666"


d.

The critical value for "\alpha = 0.05, df=n-1=22" degrees of freedom is "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}=2.073873."

"E=t_c\times\dfrac{s}{\sqrt{n}}=2.073873\times\dfrac{3.2}{\sqrt{23}}=1.3838"



e.

The critical value for "\alpha = 0.01, df=n-1=24" degrees of freedom is "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}=2.79694."

"E=t_c\times\dfrac{s}{\sqrt{n}}=2.79694\times\dfrac{2.6}{\sqrt{25}}=1.4544"



2.

a.


"\mu=28\pm2.3187"


"(25.6813, 30.3187)"

b.


"\mu=50\pm2.2380"


"(47.7620, 52.2380)"

c.


"\mu=68.2\pm0.9666"


"(67.2334, 69.1666)"

d.


"\mu=80.6\pm1.3838"


"(79.2162, 81.9838)"

e.


"\mu=92.8\pm1.4544"



"(91.3456, 94.2544)"